An amount of money was invested in an account which earned interest at a rate compounded continuously. After 11 years, the investment doubled. Determine the interest rate used at the time of investment.
Answer by stanbon(75887)

An amount of money was invested in an account which earned interest at a rate compounded continuously. After 11 years, the investment doubled. Determine the interest rate used at the time of investment.
------
Formula: A(t) = Ao*e^(rt)
Ao is the amount invested.
A(t) is the value after t years.
---
When A(t) = 2*Ao the investment has doubled.
--------
Solve: e^(rt) = 2 for "r" when t = 11.
e^(11r) = 2
-------
Take the natural log:
11r = ln(2)
r = 0.0630
rate = 6.3%
==============
Cheers,
Stan H.
==============